I can not figure this out. Please help me, I just can not get word problems correct. Thanks so much for your help. I need this help by tomorrow please.
In 1991, the life expectancy of males in a certain country 68.9 years. In 1997 it was 71.7 years. Let E represent the life expectancy in year t and let t represent the number of years since 1991.
The linear function that fits the data is. E(t)= _t + _ ? Round answer to the nearest tenth.
Also, Use the function to predict the life expectancy of males in 2009.

Answer by stanbon(75887)

In 1991, the life expectancy of males in a certain country 68.9 years. In 1997 it was 71.7 years. Let E represent the life expectancy in year t and let t represent the number of years since 1991.
The linear function that fits the data is. E(t)= _t + _ ? Round answer to the nearest tenth.
Also, Use the function to predict the life expectancy of males in 2009.
-------------------
I'm sure you know how to find the equation of a line if you have two points.
Well, in this problem you are given two points: (0,68.9) and (6,71.7)
slope = (71.7-68.9) / (6-0) = 0.466667
Form is E(t) = slope(years since 1991) + b
Solve for "b"
68.9 = 0.466667*0 + b
b= 68.9
--------------
EQUATION:
E(t) = 0.466667t + 68.9
===========================
Use the function to predict the life expectancy of males in 2009.
E(2009-1991) = 0.46666667*18 + 68.9
E(18) = 77.3 yrs.
=====================
Cheers,
Stan H.